OpenStudy (anonymous):

A man has two more nickels than dimes. He has $1.15 in all. How many coins of each kind does he have? Which equation can be used to solve the problem? 5x + 10(x + 2) = 115 5(x + 2) + 10x = 115 x + x + 2 = 115

OpenStudy (anonymous):

5(x + 2) + 10x = 115

OpenStudy (anonymous):

5x + 10(x + 2) = 115

OpenStudy (anonymous):

5(x+2+2x)=115